40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 21:05:23 2022, yuta, Update, BHD, AS path alignment, removing a lot of green stuff 7x
    Reply  Wed May 4 13:09:42 2022, Anchal, Update, BHD, IMC table balanced again, IMC is locking, YARM is locking 
       Reply  Wed May 4 15:08:59 2022, Tega, Update, BHD, XARM alignment to get flashing IFO_XY_ARMS_Flashing.png
          Reply  Wed May 4 18:52:43 2022, Tega, Update, BHD, POP beam too high at POP SM4 on the ITMX chamber snnp_pr2_pr3_itmy_aligment.png
             Reply  Thu May 5 14:46:22 2022, Anchal, Update, BHD, POP beam height lowered, POP_SM4 raised 
                Reply  Thu May 5 17:05:31 2022, Tega, Update, BHD, IMC & X/Y-arm alignment IFO_aligment_Y_locking_on.png
                Reply  Mon May 9 18:43:03 2022, Anchal, Update, BHD, ITMX table layout corrected 
                   Reply  Mon May 9 18:49:05 2022, Tega, Update, BHD, Relocate green TRX and TRY PDs/cams/optics from PSL table to BS table IMG_20220509_184439943.jpgIMG_20220509_184154722.jpg
                      Reply  Tue May 10 15:46:38 2022, JC, Update, BHD, Relocate green TRX and TRY components from PSL table to BS table IMG_0741.jpegIMG_0753.jpeg
                         Reply  Tue May 10 18:25:43 2022, Paco, Update, BHD, Green Transmission path 
                   Reply  Tue May 10 16:31:14 2022, Tega, Update, BHD, ITMX optlev return beam steered to QPD 
       Reply  Fri May 6 13:48:34 2022, Anchal, Update, BHD, WFS loop instability fixed 
    Reply  Thu May 12 19:55:01 2022, Tega, Update, BHD, AS path alignment AS.JPGOMCchamber.JPG
       Reply  Thu May 12 20:11:18 2022, Anchal, Update, BHD, BHDBS Output beams steered out to ITMY table signal-2022-05-12-201844.jpeg
          Reply  Thu May 12 20:24:29 2022, yuta, Update, BHD, POX and POY investigation POY.JPG
       Reply  Mon May 16 13:22:59 2022, yuta, Update, BHD, REFL and AS paths aligned at AP table REFLPOP.JPGPOPAS.JPG
          Reply  Mon May 16 19:14:17 2022, Anchal, Update, BHD, Camera set on AS path and BHDBS output path 
             Reply  Tue May 17 18:43:38 2022, Anchal, Update, BHD, Placed SRM in ITMY Chamber 
                Reply  Wed May 18 08:30:29 2022, Paco, Update, BHD, SRM OpLev 
                   Reply  Thu May 19 08:51:40 2022, Paco, Update, BHD, SRM OpLev 
                Reply  Wed May 18 09:02:52 2022, Anchal, Update, BHD, WFS1 PD centered 
                Reply  Wed May 18 17:23:15 2022, Anchal, Update, BHD, Placed PRM in BS Chamber 
                   Reply  Thu May 19 18:56:08 2022, yuta, Update, BHD, Current OSEM sensor values with all the suspensions aligned OSEMS.txt
                   Reply  Thu May 19 19:05:59 2022, Anchal, Update, BHD, BS Chamber all work finished, BHD path setup 
                      Reply  Fri May 20 20:03:48 2022, Paco, Update, BHD, ITMY chamber work finished - LO and AS overlapped Screenshot_2022-05-20_20-38-25.pngOSEMs.txt
                         Reply  Mon May 23 13:16:59 2022, Paco, Update, BHD, End of vent - checks OSEMs20220523.txtScreenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.pngPR3position.JPG
                            Reply  Tue May 24 15:21:13 2022, Anchal, Update, BHD, Freeswing tests of new SOS started 
Message ID: 16849     Entry time: Thu May 12 20:11:18 2022     In reply to: 16848     Reply to this: 16850
Author: Anchal 
Type: Update 
Category: BHD 
Subject: BHDBS Output beams steered out to ITMY table 

I successfully steered out the two output beams from BHD BS to ITMY table today. This required significant changes on the table, but I was able to bring back the table to balance coarsely and then recover YARM flashing with fine tuning of ITMY.

  • The counterweights were kept at the North end of the table which was in way of one of the output beams of BHD.
  • So I saved the level meter positions in my head and removed those counterweights.
  • I also needed to remove the cable post for ITMY and SRM that was in the center of the table.
  • I installed a new cable post which is just for SRM and is behind AS2. ITMY's cable post is next to it on the other edge of the table. This is to ensure that BHD board can come in later without disturbing existing layout.
  • I got 3 Y1-45P and 1 Y1-0 mirror. The Y1-0 mirror was not installed on a mount, so I removed an older optic which was unlabeled and put this on it's mount.
  • Note that I noticed that some light (significant enough to be visible on my card) is leaking out of the 45P mirrors. We need to make sure we aren't loosing too much power due to this.
  • Both beams are steered through the center of the window, they are separating outside and not clipping on any of the existing optics outside. (See attachment 1, the red beam in the center is the ITMY oplev input beam and the two IR beams are the outputs from BHD BS).
  • Also note that I didn't find any LO beam while doing this work. I only used AS beam to align the path.
  • I centered the ITMY oplev at the end.

Next steps:

  • LO path needs to be tuned up and cleared off again. We need to match the beams on BHD BS as well.
  • Setup steering mirrors and photodiodes on the outside table on ITMY.
Attachment 1: signal-2022-05-12-201844.jpeg  249 kB  Uploaded Thu May 12 21:21:34 2022  | Hide | Hide all
signal-2022-05-12-201844.jpeg
ELOG V3.1.3-